### v1

# Exploring the Relationship between Analogy Identification and Sentence Structure Encoding in Large Language Models

Abstract
Identifying analogies plays a pivotal role in human cognition and language proficiency. In the last decade, there has been extensive research on word analogies in the form of “A is to B as C is to D.” However, there is a growing interest in analogies that involve longer text, such as sentences and collections of sentences, which convey analogous meanings. While the current NLP research community evaluates the ability of Large Language Models (LLMs) to identify such analogies, the underlying reasons behind these abilities warrant deeper investigation. Furthermore, the capability of LLMs to encode both syntactic and semantic structures of language within their embeddings has garnered significant attention with the surge in their utilization. In this work, we examine the relationship between the abilities of multiple LLMs to identify sentence analogies, and their capacity to encode syntactic and semantic structures. Through our analysis, we find that analogy identification ability of LLMs is positively correlated with their ability to encode syntactic and semantic structures of sentences. Specifically, we find that the LLMs which capture syntactic structures better, also have higher abilities in identifying sentence analogies.
1 Introduction
Analogies facilitate the transfer of meaning and knowledge from one domain to another. Making and identifying analogies is a central tenet in human cognition Hofstadter (2001); Holyoak et al. (2001) and is aided by humans’ ability to process the structure of language. In the domain of NLP, several types of textual analogies are identified, such as word analogies Yuan et al. (2023); Gladkova et al. (2016); Gao et al. (2014), proportional word analogies Chen et al. (2022); Ushio et al. (2021); Szymanski (2017); Drozd et al. (2016), sentence-analogies Afantenos et al. (2021); Zhu and de Melo (2020); Wang and Lepage (2020) and more recently analogies between procedural/long text Sultan and Shahaf (2022).
Figure 1: This pipeline details the process of quantifying the LLMs abilities to capture sentence structure via SyntScore and SemScore scores for a given sentence. In this work, we apply this process to a dataset of 100K sentences. The dataset is divided into 0.8 for training the structure probe and 0.1 for testing.
Despite the existence of several established benchmarks (e.g., SuperGLue and GLUE) which evaluate the abilities of LLMs extrinsically, Wijesiriwardene et al. (2023) propose a more challenging intrinsic benchmark that focuses on LLMs’ ability to identify analogies across a range of complexities. Researchers argue that language understanding necessitates humans to grasp the inherent structure of language, including parse trees and logical formulas Chomsky (2002); Schank (1972). Furthermore, the identification of analogies relies on the utilization of implicit relational knowledge embedded within the relational structure of language Gentner (1983). Consequently, the present study investigates the potential relationship between LLMs’ ability to identify analogies and their aptitude for encoding sentence structures.
This work looks explicitly at sentence-level analogies  [^1] An analogy in the context of our manuscript always means sentence-analogy., which is a clear starting point for this work. We also look at LLMs’ abilities to capture both syntactic and semantic structures of sentences. Specifically, our main contribution is an empirical analysis of the correlation between the analogy identification ability and sentence structure encoding abilities of LLMs. Additionally, we extend the sentence structure probing techniques introduced by Hewitt and Manning (2019) (which only supports BERT and ELMo) to further work with encoder-decoder-based LLMs and LLMs that use two transformer architectures. Finally, we utilize the structure probing technique originally used for syntactic structure probing in the novel context of semantic structure probing.
2 Related Work
Assessing the ability of Neural Networks (NN) to encode syntactic and semantic structures of language is well examined in NLP Nivre et al. (2007); Manning and Schutze (1999); Parsing (2009). Everaert et al. (2015) emphasize that the meaning of sentences is inferred by the hierarchical structures provided by syntactic and semantic properties of language.
Syntactic parsing aims to derive the syntactic dependencies in a sentence, such as subjects, objects, quantifiers, determiners and other similar elements. Early probing tasks Adi et al. (2016); Shi et al. (2016) tried to identify NNs’ abilities to capture syntactic structures by classifying sentences with single and plural subjects. Later, Conneau et al. (2018) showed that NNs could capture the maximal parse tree depth. The structure probing technique used and extended in this work Hewitt and Manning (2019) is related but distinct due to its ability to implicitly capture the parse tree structures through simple distance measures between the vector representations of the words.
Compared to syntactic parsing, the NLP communities’ interest in semantic parsing is growing. Semantic parsing maps natural language sentences to a complete, formal meaning representation. Semantic parsing is achieved via combining the Semantic Role Labelling (SRL) approaches with syntactic dependency parsing Hajic et al. (2009); Surdeanu et al. (2008) and more recently via semantic dependency parsing Oepen et al. (2014, 2015). This work uses the semantic dependency parsing approach based on mean field variational inference (MFVI) augmented with character and lemma level embeddings introduced by Wang et al. (2019).
3 Approach
Our approach to exploring the relationship between analogy identification and sentence structure encoding in LLMs is detailed in the following three subsections. We explain the dataset used, in Section 3.1, the analogy identification abilities of LLMs in Section 3.2 and the sentence structure encoding abilities of LLMs in Section 3.3.
3.1 Dataset
We experiment on a dataset of 100K English sentences. Specifically, the dataset used in this work is randomly picked from the sentence corpus of levels three, four and five of the analogy taxonomy introduced in Wijesiriwardene et al. (2023). The composition of the dataset is presented in Table 1 (duplicates removed).
Analogy Taxo. Level Datasets # Sentences
Level Three Random deletion/masking/reorder 69,111
Level Four Negation 1,245
Level Five Entailment 29,644
Total # Sentences 100,000
Table 1: Dataset statistics.
3.2 Large Language Models and their Ability to Capture Sentence Analogies
We experiment on the eight language models used in a study by Wijesiriwardene et al. (2023) namely, BERT Devlin et al. (2018), RoBERTa Liu et al. (2019), ALBERT Lan et al. (2019), LinkBERT Yasunaga et al. (2022), SpanBERT Joshi et al. (2020) and XLNet Yang et al. (2019) which are encoder-based LLMs, T5 Raffel et al. (2020), an encoder-decoder-based LLM and ELECTRA Clark et al. (2020), an LLM based on two transformer architectures. We refer readers to cited publications for details on the LLMs used in this work.
Wijesiriwardene et al. (2023) introduced a taxonomy of analogies starting from less complex word-level analogies to more complex paragraph-level analogies and evaluated how each LLM performs on identifying analogies at each level of the taxonomy. An analogy is a pair of lexical items that are identified to hold a similar meaning to each other. Therefore the distance between a pair of analogous lexical items in the vector space should be smaller. The same authors identify Mahalanobis Distance (MD) Mahalanobis (1936) to be a better measurement of the distance between two analogous sentences in the vector space. Therefore in this work, the ability of each LLM to identify sentence analogies is represented by the mean MD calculated for the sentence-level datasets (levels 3, 4 and 5) present in the analogy taxonomy. These mean values are calculated based on the reported values by Wijesiriwardene et al. (2023).
3.3 Large Language Models and their Ability to Capture Sentence Structures
Hewitt and Manning (2019) introduced a probing approach to evaluate whether syntax trees (sentence structures) are encoded in Language Models’ (LMs’) vector geometry. The probing model is trained on train/dev/test splits of the Penn Treebank Marcus et al. (1993) and tested on both BERT Devlin et al. (2018) and ELMo Peters et al. (2018). An LM’s ability to capture sentence structure is quantified by its ability to correctly encode the gold parse tree (provided in the Penn Treebank dataset) within its embeddings for a given sentence.
The authors introduce a path distance metric and a path depth metric for evaluation. The distance metric captures the path length between each pair of words measured by Undirected Unlabeled Attachment Score (UUAS) and average Spearman correlation of true to predicted distances (DSpr). The depth metric evaluates the model’s ability to identify the sentence’s root, measured as root accuracy percentage. Additionally, the depth metric also evaluates the ability of the model to recreate the word order based on their depth in the parse tree identified as Norm Spearman (NSpr.) [^2] We do not use NSpr. in this work. We refer the readers to Hewitt and Manning (2019) for further details on the technique and evaluation metrics.
4 Experimental Setup
Exploring the relationship between analogy identification and sentence structure encoding abilities of LLMs requires a representative score to quantify (i) analogy identification ability (AnalogyScore), (ii) semantic structure identification ability (SemScore), and (iii) syntactic structure identification ability (SyntScore) of each LLM.
We obtain AnalogyScore by calculating the means of reported MD measures obtained for each sentence-level dataset in Wijesiriwardene et al. (2023).
To obtain the SemScore (see Figure 1), we first parse all the sentences in our dataset using the MFVI approach Wang et al. (2019). The resulting semantically parsed sentences (in CoNLL-U format) [^3] https://universaldependencies.org/format.html and the LLM embeddings of the original sentences are sent to the structure probing technique Hewitt and Manning (2019). The structure probe is trained on 80K sentences from the dataset and results, DSpr and UUAS representing parse distance and root accuracy (RootAcc) representing parse depth are reported on the test split with 10K sentences. Finally, the SemScore is computed as a combined score by taking the mean of the z-score normalizations of these three measures $Z_{DSpr},Z_{UUAS},Z_{RootAcc}$ (see Table 2).
$\texttt{SemScore}=\frac{1}{3}(Z_{DSpr}+Z_{UUAS}+Z_{RootAcc})$
To obtain the SyntScore (see Figure 1), we follow the same steps but parse the sentences syntactically.
Finally, we calculate the Spearman’s rank correlation (SRC) and Kendall’s rank correlation (KRC) between AnalogyScore and SyntScore, as well as AnalogyScore and SemScore.
4.1 Implementation Details
When extending the structure probing technique by Hewitt and Manning (2019) to facilitate additional LLMs, we use the HuggingFace implementation [^4] https://huggingface.co/models of the LLMs. For semantic parsing, we use the trained mean field variational inference (MFVI) model augmented with character and lemma-level embeddings provided by the SuPar [^5] https://github.com/yzhangcs/parser package. For syntactic parsing of the sentences we employ Stanford CoNLL-U dependency parser [^6] https://nlp.stanford.edu/software/nndep.html.
Model Original Scores Normalized Scores
Syntactic Semantic Syntactic Semantic
Distance Depth Distance Depth Distance Depth Distance Depth
DSpr UUAS RootAcc DSpr UUAS RootAcc $Z_{DSpr}$ $Z_{UUAS}$ $Z_{RootAccu}$ $Z_{DSpr}$ $Z_{UUAS}$ $Z_{RootAccu}$
ALBERT 0.59 0.46 0.35 0.38 0.13 0.19 -1.56 -2.30 -2.58 0.39 -1.30 0.36
BERT 0.73 0.72 0.74 0.38 0.16 0.17 0.87 0.62 0.56 0.39 -0.03 0.07
Electra 0.70 0.76 0.75 0.38 0.14 0.15 0.34 1.01 0.63 0.39 -0.73 -0.28
LinkBERT 0.70 0.68 0.69 0.38 0.15 0.05 0.33 0.18 0.15 0.37 -0.27 -1.79
RoBERTa 0.74 0.74 0.73 0.38 0.16 0.29 1.06 0.77 0.49 0.37 0.25 1.89
SpanBERT 0.74 0.72 0.74 0.38 0.14 0.20 1.06 0.56 0.55 0.37 -0.97 0.54
T5 0.63 0.64 0.71 0.37 0.19 0.17 -0.79 -0.31 0.28 -2.65 1.64 0.05
XLNet 0.60 0.62 0.66 0.38 0.18 0.11 -1.31 -0.53 -0.08 0.37 1.42 -0.83
Table 2: DSpr, UUAS measures indicating Parse Distance (Distance) and RootAcc measure indicating Parse Depth (Depth). Original Scores denote original output values of the structure probe technique and Normalized Scores are z-score normalized. Higher values indicate a stronger ability of the LLMs to capture sentence structures.
Model AnalogyScore SyntScore SemScore
Score Rank Score Rank Score Rank
ALBERT 0.645 7 -2.14 8 -0.19 5
BERT 0.505 3 0.68 3 0.14 3
Electra 0.516 4 0.66 4 -0.21 6
LinkBERT 0.608 6 0.22 5 -0.56 8
RoBERTa 0.458 1 0.78 1 0.84 1
SpanBERT 0.461 2 0.72 2 -0.02 4
T5 0.524 5 -0.27 6 -0.32 7
XLNet 0.747 8 -0.64 7 0.32 2
Table 3: The values for AnalogyScore, SyntScore and SemScore and their corresponding rank values. AnalogyScore ranges between [0,1], 0 being the best. For SyntScore and SemScore higher the values better the ability of LLMs to capture sentence structure.
5 Results
5.1 Semantic and Syntactic Structure Encoding Abilities of LLMs
We tabulate the structure probing results in original metrics (Table 2) and the performance of each LLM in identifying sentence analogies and capturing the semantic and syntactic structures (Table 3). It is interesting to note that RoBERTa, the best-performing LLM for analogy identification (AnalogyScore $=0.458$), holds the highest SyntScore and SemScore. XLNet is the lowest-performing model for analogy identification as well as syntactic structure identification. Yet it performs second-best in semantic structure identification. SpanBERT ranks second in both analogy identification and syntactic structure identification but holds the median SemScore.
5.2 Analogy Identification and Syntactic Structure Encoding Abilities of LLMs
We use SRC and KRC values to analyze the correlation between LLMs’ ability to identify sentence analogies denoted by AnalogyScore and LLMs’ ability to encode syntactic structures of sentences denoted by SyntScore. Both correlation measures show a positive correlation between AnalogyScore and SyntScore. Specifically, the SRC between AnalogyScore and SyntScore is 0.95 ($p<0.001$). The KRC between AnalogyScore and SyntScore is 0.86 ($p=0.002$).
5.3 Analogy Identification and Semantic Structure Encoding abilities of LLMs
Similar to the previous section, we compute the SRC and KRC values to asses the correlations between AnalogyScore and SemScore. We see that both correlations are positive with SRC of 0.33 ($p=0.42$) and KRC of 0.28 ($p=0.40$) between AnalogyScore and SemScore.
6 Limitations
Even though Abstract Meaning Representation (AMR) Banarescu et al. (2013) is a popular and widely used technique to parse sentence meanings, in this work, we use MFVI, a semantic parsing approach introduced by Wang et al. (2019) because of the limitations posed by the structure probing technique used Hewitt and Manning (2019). This technique requires the mapped LLM embeddings and semantic dependency parsed sentences to be the same size. However, as it is known, AMRs abstract away from the syntactic idiosyncrasies of the language and overlook certain auxiliary words from the parse results, limiting its use in this work.
The present study employs a semantic parsing technique reported to exhibit a high accuracy level of 94% (Wang, 2019). However, it is important to note that for the purposes of our investigation, we make the assumption that the semantically parsed sentences generated by this particular method are entirely accurate, thereby employing them as the gold standard data. It is worth mentioning that this choice may introduce some degree of bias into our examination of the semantic structure probing.
We obtain sentence-analogy pairs provided by Wijesiriwardene et al. (2023) [^7] We thank the authors of Wijesiriwardene et al. (2023) for providing access to the datasets. and split the pairs to obtain single sentences. However, it should be noted that the subset of sentences employed in this work does not exhibit a balanced distribution across the three levels of the taxonomy.
7 Conclusion and Future Avenues
This work explores the correlation between LLMs’ ability to identify analogies and encode sentence structures in their embeddings. Through detailed experiments, we show that the analogy identification ability of LLMs is positively correlated with their ability to encode syntactic and semantic structures. Particularly, LLMs that better capture syntactic structures have a higher correlation to analogy identification. As a next step, we would like to assess the possibility of extending this approach in the context of explainability of LLMs.